OpenStudy (anonymous):

another where did i go wrong:

OpenStudy (anonymous):

Before posting and solving the question you went wrong?? This is strange..

OpenStudy (anonymous):

*go

OpenStudy (angela210793):

he went wrong while thinking how to solve it @waterineyes :P

OpenStudy (lgbasallote):

ahh glad to see the old gang together again

OpenStudy (anonymous):

look now you draw a huge crowd lol

OpenStudy (angela210793):

@waterineyes excuse me?

OpenStudy (anonymous):

theres no pressure tho be cool

OpenStudy (anonymous):

\[10^{2} \times 10^{\frac{ 1 }{ 2}} \times 10^{\frac{ -1 }{ 3 }} \times10^{-6} \times 10^{x}\times10^{\frac{ -1 }{ 2}x}\]

OpenStudy (anonymous):

people please, i am not providing entertainment just looking for answers

OpenStudy (anonymous):

Try it once..

OpenStudy (anonymous):

yes i'll

OpenStudy (anonymous):

wait angela

OpenStudy (angela210793):

there was a ?=question mark at the end if u didn't notice @waterineyes

OpenStudy (anonymous):

oh okay

OpenStudy (angela210793):

I thought so..

OpenStudy (anonymous):

10^2 x 10^1/2 x10^-1/3 x10^-6 = \[\frac{ 10^{\frac{ 5 }{ 2 }} }{ 10^{\frac{ -19 }{ 3 }} }\] 10^x X 10^-1/2x =\[\frac{ 10^{x} }{ 10^{\frac{ 1 }{ 2}x} }\]

OpenStudy (anonymous):

good so farr?

OpenStudy (anonymous):

See it will be easier if you add all the powers first...

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

\[\large x^a \times x^b \times x^c \times x^d \times x^e \implies x^{a + b+c+d+e}\]

OpenStudy (anonymous):

No need to write the base just add all the powers first...

OpenStudy (anonymous):

2+1/1-1/3-6= \[\frac{ 6+3-2-6 }{ 6 } = \frac{ 1 }{ 6}\]

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

\[2 + \frac{1}{2} - \frac{1}{3} - 6 = ??\]

OpenStudy (anonymous):

Take your time..

OpenStudy (anonymous):

common denominator: 6 \[\frac{ 6+3-2-6 }{ 6}\]

OpenStudy (anonymous):

isn't it?

OpenStudy (anonymous):

Check for last part that is 6..

OpenStudy (anonymous):

ah

OpenStudy (anonymous):

-36

OpenStudy (anonymous):

make the common denominator for 6 also..

OpenStudy (anonymous):

Yep.. -36

OpenStudy (anonymous):

Now solve it..

OpenStudy (anonymous):

9-38/6= 29/6

OpenStudy (anonymous):

-29

OpenStudy (anonymous):

Yep...

OpenStudy (anonymous):

Now: \[x - \frac{1}{2}x = ??\]

OpenStudy (anonymous):

1/2x

OpenStudy (anonymous):

So what do you get now ??

OpenStudy (anonymous):

Write it in powers of 10 ..

OpenStudy (anonymous):

10^-29/6+1/2x

OpenStudy (anonymous):

but thats wrong..

OpenStudy (anonymous):

my book says 10^-23/6+3/2x

OpenStudy (anonymous):

More clearly: \[\large 10^{{\frac{1}{2}}x + \frac{29}{6}}\]

OpenStudy (anonymous):

Then we have to check it again..

OpenStudy (anonymous):

unless the book is wrong..

OpenStudy (anonymous):

x - 1/2x = 1/2 x It can't be 3/2x unless x + 1/2x is there..

OpenStudy (anonymous):

wait i didn't see the last power right its 10^1/2x

OpenStudy (anonymous):

so the 3/2x is accounted for

OpenStudy (anonymous):

Sorry I did mistake of minus here: \[\large 10^{{\frac{1}{2}}x - \frac{29}{6}}\]

OpenStudy (anonymous):

In that case it has become now: \[\large 10^{{\frac{3}{2}}x - \frac{29}{6}}\]

OpenStudy (anonymous):

but the fraction is still -23/6

OpenStudy (anonymous):

We have calculated it right.. May be misprint in book..

OpenStudy (anonymous):

Let me check the powers by which we will get -23/6..

OpenStudy (anonymous):

i guess book misprint :) thank you

OpenStudy (anonymous):

Is your question right ??

OpenStudy (anonymous):

yes, i corrected the last power

OpenStudy (anonymous):

No no except that..

OpenStudy (anonymous):

everything else is right

OpenStudy (anonymous):

If you multiply one 10 with your question then you will get -23/6

OpenStudy (anonymous):

-29/6 + 1 = -23/6 Ha ha ha Nevermind.. Book misprint..

OpenStudy (anonymous):

yep